OpenStudy (anonymous):

PLEASE i need help, ill fan and medal<3 The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (rhystic):

the area of the prism is defined as A = 2*b+p*h where b is the area of the base p is the perimeter and h is the height. This is because regardless of the shape of the prism the prism will have two sides that match the base (the bottom and the top) this is where the 2*b comes in. all the stuff in the middle regardless of the shape will match the shape of the base meaning if you just take the perimeter and multiply by the height you cover the area in between. so you have 2*55+30*7 = 110+210 = 320